# ABC-SN: Attention Based Classifier for Supernova Spectra

### v1

@@ 1 Introduction
SNe Ib, Ic, and II are core-collapse SNe (Janka et al., 2012; Modjaz et al., 2019) with progenitor stars of mass of (Woosley & Weaver, 1986; Woosley & Janka, 2005; Couch, 2017). On the other hand, SNe Ia are the result of thermonuclear explosions (Röpke, 2007) produced when a white dwarf (WD) star is accreting matter from a nearby star (Mazzali et al., 2007) beyond the “Chandrasekhar limit” of (although in practice both sub- and super-Chandrasekhar progenitors have been identified, Andrew Howell et al. 2006, e.g.,; Hicken et al. 2007, e.g.,, see for example Woosley & Kasen 2011 for a theoretical approach). Carbon fusion occurs quickly and explosively, producing the SNe that are observed. The intrinsic brightness of SNe Ia can be calculated from their temporal evolution and color, and they are therefore known as “standard(izable) candles” and cosmologists use SNe Ia to study dark energy and the expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999).
@@
Additional explosion methods (e.g.,, pair instability Woosley, 2017; Kasen et al., 2011), and several more types and subtypes of Super-Luminous Supernovae exist, but are not addressed in this work; see Moriya et al. (2018) for a review.
@@
Altogether, this brief introduction only scratches the surface of a complex, and, apparently, ever-growing taxonomy. Typing SN reflects sometimes completely different progenitor or explosion mechanisms, sometimes nuanced differences in a continuum of parameters such as how much material was stripped before the explosion, mass, metallicity of the progenitor environment, etc. The reader should keep this in mind and throughout this report we discuss how this may impact the result of automated classifiers.
@@
While photometric classification methods (e.g., Boone, 2019; Möller & de Boissière, 2020; Qu et al., 2021; Shah et al., 2025) have made significant advances motivated by the upcoming LSST which will discover too many and too faint SNe for comprehensive spectroscopic classification, and multimodal models are starting to appear (Zhang et al., 2024), classifying SNe spectroscopically remains the most common way to determine most subtypes of SNe.
@@
Currently, the most prevalent method for the classification of SN subtypes is Supernova Identification (Blondin & Tonry, 2007, SNID), which performs a cross-correlation with an extensive library of labeled SNe and template spectra. Though this approach has been a mainstay in the field due to its reliability and interpretability, the oncoming data deluge from the LSST presents a scalability challenge that SNID may struggle to meet efficiently. As SN science evolves, template matching methods require continuous, careful reclassification. Meanwhile, machine learning models like SNIascore (Fremling et al., 2021), which have shown proficiency in classifying Type Ia SNe with the low-resolution spectral data from the SEDMachine (Blagorodnova et al., 2018, SEDM), have yet to demonstrate the same adaptability for a wider variety of SN subtypes.
@@
This paper introduces ABC-SN the ‘Attention-Based Classifier for Supernovae’, a novel approach to SN classification that leverages the multi-head attention layer in the transformer neural network architecture (Vaswani et al., 2017), a model that has significantly impacted the field of natural language processing and was subsequently adapted for innumerable applications from computer vision (Dosovitskiy et al., 2020) to time series forecasting (Zhou et al., 2021). Like the complex web of syntax and semantics in human languages, SN spectra contain intricate patterns — spectral features — that signal the presence of specific elements, stellar structure patterns, and determine a SN’s subtype. We posit that the transformer’s ability to parse these complex patterns in language could be directly applicable to spectral classification. Applications of transformer-derived architectures to spectroscopy outside of astrophysics include (Cai et al., 2022) and a number of models for application on hyperspectral imaging (e.g., He et al., 2021). In astrophysics, see for example Zhang et al. (2024) and Koblischke & Bovy (2024).
@@ 2 Retraining DASH
A dataset of SN spectra poses a unique complication: the same SN may be observed multiple times at different epochs in the lifetime of the transient. For instance, our dataset is comprised of spectra from objects, where 144 SNe contribute 10 or more spectra each. While SNe do evolve, key signatures of the cosmic explosion will persist over time. Thus, all spectra from the same SN are highly correlated; they are typically more self-similar than similar to spectra of other SNe of the same subtype.
@@
Then we curtail overfitting by implementing an early stopping protocol which halts training if model performance doesn’t increase. Specifically, if test set F1-score does not increase by 0.005 for 10 pseudo-epochs, we stop training. Also, if training set F1-score exceeds the test set by 0.10 for 10 pseudo-epochs, we halt training. This satisfies our second goal (Figure 1, middle).
@@ 3.3 Culling, Standardizing and Validating
Each spectrum is defined from 2,500Å to 10,000Å, but we set each spectrum’s flux value to 0 at wavelengths shorter than 4,500Å and longer than 7,000Å to enforce uniformity. While this is a restricted wavelength range it served to ensure our data was maximally homogeneous and avoid working with missing values. We also normalize each spectrum to be mean 0 and standard deviation 1, not including the 0-padding. At this point, we remove any spectra that have either a very small (0.1) or very large (100) dynamic range (the difference between a spectrum’s minimum value and maximum value).
@@
Due to the small number of SNe in some classes, we did not split our data into a training-validation-test set. Thus, our model hyperparameters are selected based on test set performance.
@@ 6 Conclusion and Future Work
ABC-SN is an open source tool222we will add a GitHub link here and also as astro software library link, we provide a pretrained model ready for application, together with tools for preprocessing spectra, and we provide the code base and hope the astrophysical community will continue to refine ABC-SN for additional applications.
@@ Appendix A ABC-SN Training hyperparameters and data preparation choices
Spectrum shifting | Shift spectrum by wavelength bins drawn from a uniform (integer) distribution in this range. | |
@@
Minimum | RLRP will not reduce below this threshold. |

### v2

@@ 1 Introduction
SNe Ib, Ic, and II are core-collapse SNe (Janka et al., 2012; Modjaz et al., 2019) with progenitor stars of mass of (Woosley & Weaver, 1986; Woosley & Janka, 2005; Couch, 2017). On the other hand, SNe Ia are the result of thermonuclear explosions (Röpke, 2007) produced when a white dwarf (WD) star is accreting matter from a nearby star (Mazzali et al., 2007) beyond the “Chandrasekhar limit” of (although in practice both sub- and super-Chandrasekhar progenitors have been identified, Andrew Howell et al. 2006, e.g.,; Hicken et al. 2007, e.g.,, see for example Woosley & Kasen 2011 for a theoretical approach). Carbon fusion occurs quickly and explosively, producing the SNe that are observed. The intrinsic brightness of SNe Ia can be calculated from their temporal evolution and color, and they are therefore known as “standard(izable) candles”. Cosmologists use SNe Ia to study dark energy and the expansion of the Universe (Riess et al., 1998; Perlmutter et al., 1999).
@@
Additional explosion methods (e.g., pair instability Woosley, 2017; Kasen et al., 2011), and several more types and subtypes of Super-Luminous Supernovae exist, but are not addressed in this work; see Moriya et al. (2018) for a review.
@@
Altogether, this brief introduction only scratches the surface of a complex and, apparently, ever-growing taxonomy. Typing SN reflects sometimes completely different progenitor or explosion mechanisms, sometimes nuanced differences in a continuum of parameters such as how much material was stripped before the explosion, mass, metallicity of the progenitor environment, etc. The reader should keep this in mind and throughout this report we discuss how this may impact the result of automated classifiers.
@@
While photometric classification methods (e.g., Boone, 2019; Möller & de Boissiére, 2020; Qu et al., 2021; Shah et al., 2025) have made significant advances motivated by the upcoming LSST which will discover too many and too faint SNe for comprehensive spectroscopic classification, and with multi-modal models starting to appear (Zhang et al., 2024), classifying SNe spectroscopically remains the only way to determine most subtypes of SNe. Photometric classifiers can reach high accuracy in separating broad classes, but none of the photometric classifier can differentiate the stripped envelope supernova subtypes (Khakpash et al., 2024): SN IIb, Ib, Ic, Ic-broad, Ibn etc., which are usually classified together as “SN Ibc”. Currently, the most prevalent methods for the classification of SN subtypes are based on template matching: Supernova Identification (Blondin & Tonry, 2007, SNID) and SUPERFIT (Howell et al., 2005, recently automated in Python as Next Generation SuperFit NGSF, Goldwasser et al. 2022) or GEneric cLAssification TOol (Harutyunyan et al., 2008, GELATO).
@@
SNID types new SNe with an extensive library of labeled SNe and template spectra (we will use the SNID library as our dataset and updated SNID labels as our ground truth, see section 3). Though this approach has been a mainstay in the field due to its reliability and interpretability, the oncoming data deluge from the LSST presents a scalability challenge that SNID may struggle to meet efficiently. As SN science evolves, template matching methods require continuous, careful reclassification. Meanwhile, machine learning models like SNIascore (Fremling et al., 2021) and its sibling CCSNscore (Sharma et al., 2025), which have shown proficiency in classifying Type Ia SNe and separating Hydrogen-rich from Hydrogen-poor SNe with the low-resolution spectral data from the SEDMachine (Blagorodnova et al., 2018, SEDM), have yet to demonstrate the same adaptability for a wider variety of SN subtypes and spectra sourced from different instruments.
@@
This paper introduces ABC-SN the ‘Attention-Based Classifier for Supernovae’, a novel approach to SN classification that leverages the multi-head attention layer in the transformer neural network architecture (Vaswani et al., 2017), a model that has significantly impacted the field of natural language processing and was subsequently adapted for innumerable applications from computer vision (Dosovitskiy et al., 2020) to time series forecasting (Zhou et al., 2021). Like the complex web of syntax and semantics in human languages, SN spectra contain intricate patterns — spectral features — that signal the presence of specific elements, indicate stellar structure patterns, and determine a SN’s subtype. We posit that the transformer’s ability to parse these complex patterns in language could be directly applicable to spectral classification. Applications of transformer-derived architectures to spectroscopy outside of astrophysics include (Cai et al., 2022) and a number of models for application on hyperspectral imaging (e.g., He et al., 2021). In astrophysics, see for example Zhang et al. (2024) and Koblischke & Bovy (2024).
@@ 2 Retraining DASH
A dataset of SN spectra poses a unique complication: the same SN may be observed multiple times at different epochs in the lifetime of the transient. For instance, our dataset is comprised of spectra from SNe, where 144 SNe contribute 10 or more spectra each. While SNe do evolve, key signatures of the cosmic explosion will persist over time. Thus, all spectra from the same SN are highly correlated; they are typically more self-similar than similar to spectra of other SNe of the same subtype.
@@
Then we curtail overfitting by implementing an early stopping protocol which halts training if model performance doesn’t increase. Specifically, if test set F1-score does not increase by 0.005 for 10 pseudo-epochs, we stop training. Also, if the training set F1-score exceeds the test set by 0.10 for 10 pseudo-epochs, we halt training. This satisfies our second goal (Figure 1, middle).
@@ 3.3 Culling, Standardizing and Validating
Each spectrum is defined from 2,500Å to 10,000Å, but we set each spectrum’s flux value to 0 at wavelengths shorter than 4,500Å and longer than 7,000Å to enforce uniformity. While this is a restricted wavelength range it served to ensure our data was maximally homogeneous and avoided working with missing values. We also normalize each spectrum to be mean and standard deviation , not including the 0-padding. At this point, we remove any spectra that have either a very small (0.1) or very large (100) dynamic range (the difference between a spectrum’s minimum value and maximum value).
@@
To validate the stability of our results, we trained ABC-SN five times with different random seed initializations. The performance reported here is the mean performance of the five models. Due to the severe imbalance of our dataset (see Table 1) the transferability of our result should be considered with caution for the smallest classes. There is no way to ensure that the classification power will remain as measured if the class is only populated by three (Ibn) or six (Iax, IIP) unique SNe as we simply do not have a sample large enough to assume our data is representative. Similarly, the constraint on keeping all augmented copies of the same SN in either the training or test set does not allow us to proceed with a traditional cross-validation scheme.
@@ 5 Results and Discussion
Harder yet is a comparison with CCSNscore since the model architecture is markedly different: it is a multi-modal (incorporating both lightcurves and spectra) hierarchical model, and their classes are, in some cases, defined differently than ours. Nonetheless, based on their reported performance on subtypes (Section 5.2 and Figure 6 in Sharma et al. 2025) we outperform them on all classes in common: their spectra-only classification accuracy for types (IIb, Ib, Ic, Ic-bl) is (0.15, 0.44, 0.16, 0.50) compared to our (0.70, 0.82, 0.82, 0.82), and their accuracy only improves for IIb when other data modalities (photometry) are included (0.55, 0.36, 0.14, 0.17). We caution the reader, however, that comparison between models should only be done on shared, benchmark datasets (which motivated our work in re-training DASH), so the comparisons with SNIascore and CCSNscore should be treated cautiously.
@@ 6 Conclusion and Future Work
ABC-SN is an open source tool222https://github.com/FoxFortino/ABC-SN, we provide a pretrained model ready for application333https://zenodo.org/records/16620817, together with tools for preprocessing spectra, and we provide the code base and hope the astrophysical community will continue to refine ABC-SN for additional applications.
@@ Appendix A ABC-SN Training hyperparameters and data preparation choices
| Spectrum shifting | Shift spectrum by wavelength bins drawn from a uniform (integer) | |
| distribution in this range. | ||
@@
| Minimum | RLRP will not reduce below this threshold. | |